# Self-intersection Number of Negative Curves on Fermat Surfaces

### v2

@@ Abstract.
We give an explicit formula for the self-intersection number of negative curves on Fermat surfaces. The formula offers us hints to either prove or disprove the Bounded Negativity Conjecture for the Fermat surfaces.
@@ 1. Introduction
Let $\mathbb{P}^{3}$ be the three-dimensional complex projective space, with homogeneous coordinate ring $S=\mathbb{C}[x,y,z,w]$. The Fermat surface of degree $d$, denoted as $X_{d}$, is the smooth hypersurface in $\mathbb{P}^{3}$ defined by the equation:
$x^{d}+y^{d}+z^{d}+w^{d}=0.$
The main objective in this paper is to establish the following theorem.
@@
Let $X_{d}\subset\mathbb{P}^{3}$ be the Fermat surface defined by $x^{d}+y^{d}+z^{d}+w^{d}=0$ for $d\geq 5$, and $C$ a reduced and irreducible curve on $X_{d}$ with $C^{2}<2-d<0$. Then
(1) There exists an irreducible homogeneous polynomial $g(x,y,z)$ which vanishes on $C$.
Denote by $e=\deg g$. Define two curves $E$ and $D$ in $\mathbb{P}^{2}$ by
$E:\ g=0\quad\mathrm{and}\quad D:x^{d}+y^{d}+z^{d}=0.$
(2) There exists $k\geq 2$ and $k|d$ such that
@@
$+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E),$
where $\mu_{p}(E)$ denotes the Milnor number of $E$ at $p$.
In particular, when $d$ is prime, then $k=d$ and
@@
This result is closely connected with an old folklore conjecture among experts in algebraic surface theory, commonly known as the Bounded Negativity Conjecture. We formulate this conjecture as follows.
@@
For any given smooth complex projective surface $X$, there exists a number $B_{X}\geq 0$ which depends only on $X$ such that for any reduced curve $C$ on $X$, the self-intersection number $C^{2}$ is bounded below by $-B_{X}$, i.e., $C^{2}\geq-B_{X}$.
The origins of this conjecture are unclear, but it can at least be traced back to the Italian mathematician Federigo Enriques (1871-1946), and it has been an open problem for one hundred years. For a more comprehensive overview, we refer to [3, p.1878].
@@
For any given smooth complex projective surface $X$, there exists a number $b_{X}\geq 0$ which depends only on $X$ such that for any reduced and irreducible curve $C$ on $X$, the self-intersection number $C^{2}$ is bounded below by $-b_{X}$, i.e., $C^{2}\geq-b_{X}$.
@@
Our result in Theorem 1.1 implies that under some conditions, the Bounded Negativity Conjecture is true for the Fermat surface $X_{d}$.
@@
for some constant $\lambda>0$ which may depend on $d$, where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$; or
$\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}<1,$
where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
On the other hand, the formula in Theorem 1.1 also gives us hints to construct counterexamples for the Bounded Negativity Conjecture on Fermat surfaces; see Section 5.
In the sequel, we will only investigate the self-intersection number of reduced and irreducible curves. For convenience, we denote $C$ such a curve on a given smooth projective surface.
It is evident that the Bounded Negativity Conjecture holds for $X$ with $-K_{X}\geq 0$, as can be seen by applying the adjunction formula:
@@
In particular, for $d\leq 4$, the bounded negativity of $X_{d}$ follows from this argument, since the canonical divisor of $X_{d}$ is $K_{X_{d}}=(d-4)H$ where $H$ is the hyperplane section. However, for $d\geq 5$, the canonical divisor $K_{X_{d}}$ is ample and $X_{d}$ is of general type. To the author’s knowledge, there are currently no effective methods available to deal with the Bounded Negativity of such surfaces, except for the weak Bounded Negativity Conjecture which will be quoted below.
Another natural approach to establishing the Bounded Negativity for a surface is to show that there are only finitely many negative curves on the surface. However, it is proved in [3, Thm.4.1] that for any $m>0$, there exists a smooth projective surface which admits infinitely many smooth irreducible curves of self-intersection number $-m$. For the Fermat surface $X_{d}$, we do not know whether the number of negative curves on $X_{d}$ is finite or not.
A third potential method to prove the Bounded Negativity Conjecture is to establish the boundedness of the geometric genera of reduced and irreducible curves. In fact, the weak Bounded Negativity Conjecture, proven in [10], states that on a given smooth projective surface $X$, the inequality
$C^{2}\geq-b_{X,g(C)}$
holds for any reduced and irreducible curve $C$, where $b_{X,g(C)}$ is a constant depending only on $X$ and the geometric genus $g(C)$ of $C$. Consequently, the Bounded Negativity Conjecture would follow if we could establish a bound for the genera $g(C)$. Nevertheless, such a bound does not exist for almost all surfaces. For the Fermat surface $X_{d}$, consider a generic curve $C$ in the linear series $|kH|$, then $C$ is a smooth irreducible curve on $X_{d}$ with geometric genus
$g(C)=\frac{1}{2}(K_{X_{d}}\cdot C+C^{2})+1=\frac{1}{2}\left(d(d-4)k+dk^{2}\right)+1,$
thus the genus $g(C)$ goes to infinity as $k$ tends to infinity. Therefore, the Bounded Negativity Conjecture for $X_{d}$ cannot be proved by bounding the genera of reduced and irreducible curves.
Other strategies to prove the Bounded Negativity Conjecture include investigating the non-isomorphic surjective endomorphisms or establishing dominance by other surfaces that have bounded negativity [3, Prop.2.1], exploring the birational invariance of bounded negativity and the $H$-constant [2], using the equivalence between bounded negativity and the boundedness of the denominators in the Zariski decomposition [4], discussing the weighted Bounded Negativity Conjecture [11, 8], and studying the connections between the Bounded Negativity Conjecture and other unsolved problems, such as SHGH conjecture or the bounded cohomology conjecture [1, 7]. However, despite providing invaluable insights in understanding the Bounded Negativity Conjecture, these methods either deal with only very special cases, e.g., rational surfaces or K3 surfaces, or just offer general frameworks that reduce the conjecture to other open questions. They rely on general formulas, such as the adjunction formula and Riemann-Roch formula, to establish bounds for the intersection number $C^{2}$ of a reduced and irreducible curve $C$, without using specific information about $C$, such as its equation.
Can we estimate $C^{2}$ from the equation of the curve $C$? Can we give a concrete example of a surface with the Bounded Negativity by doing explicit computation through the defining equations of the surface?
@@
Suppose $X$ and $Y$ are two smooth complex projective surfaces that are in a single family of complex projective surfaces, then the Bounded Negativity Conjecture is true for $X$ if and only if it is true for $Y$.
Assuming the conjecture above, to establish the Bounded Negativity Conjecture for surfaces in $\mathbb{P}^{3}$, it suffices to prove its validity for a specific surface. Our Theorem 1.1 is proved for this purpose.
The formulas in Theorem 1.1 show how the Bounded Negativity Conjecture is connected with the singularity theory of curves. The key term in the expression of the self-intersection number $C^{2}$ is the sum $\sum\limits_{p\in D\cap E}\mu_{p}(E)$. It is notable that this sum is taken over the intersection points of $D$ and $E$, not all the singular points of $E$.
To conclude the introduction, we give an outline of the proof of Theorem 1.1. Given a negative curve $C$ on $X_{d}$, we represent the Fermat surface $X_{d}$ as a branched Galois cover of $\mathbb{P}^{2}$:
$\rho:\ X_{d}\to\mathbb{P}^{2},\quad[x,y,z,w]\mapsto[x,y,z]$
Let $E=\rho(C)$ and $\mathscr{C}=\rho^{*}E$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$; denote by $g$ its defining equation. It is clear that $g$ can be seen as a section of the line bundle $\mathscr{O}_{X_{d}}\left((\deg g)H\right)$ which vanishes on $C$. Then we prove that $\mathscr{C}$ is a reduced curve on $X_{d}$; see Proposition 4.4. Moreover, $\mathscr{C}$ is of the following form
@@
Hence, to compute $C^{2}$, we need to calculate the sum on the right hand side. Finally, the intersection numbers can be calculated in a local way at any intersection point. Locally, the curve $\mathscr{C}$ can be represented as the curve in $(\mathbb{C}^{3},0)$ defined by
@@
The author would like to thank the organizers of International Symposium on Singularities and Applications, Dec. 9–13, 2024 in Sanya, China. The stimulating atmosphere of the conference inspired him to conceive the idea of performing computation using Puiseux series. The author was partially supported by the grant YSBR-001 and is currently supported by NSFC No. 12301100.
@@ 2. Curves on surfaces of Sebastiani-Thom type
Assume that $X\subset\mathbb{P}^{3}$ is a smooth surface defined by the following equation:
@@
where $f$ is a form of degree $d$. Such a surface is said to be of Sebastiani-Thom type. The Fermat surface is a typical example. Let $C$ be a reduced and irreducible curve on $X$.
In this section, we show that $C$ is “almost” defined by a homogeneous polynomial on $X$ in the following result.
@@
Let $X:f(x,y,z)+w^{d}=0$ be a smooth surface in $\mathbb{P}^{3}$ and $C$ an reduced and irreducible curve on $X$. Then there exists an irreducible form $g(x,y,z)$ satisfying the following properties:
  1. (i)
    The divisor on $X$ cut out by $g$ can be written as
@@
    where $C_{1}=C,C_{2},\ldots,C_{k}$ are distinct reduced and irreducible curves and $m\geq 1$ is an integer.
  2. (ii)
    The self-intersection numbers $C_{i}^{2},i=1,\ldots,k$ are all equal to $C^{2}$.
@@
It is a branched Galois covering map with Galois group $\mathbb{Z}_{d}=\mathbb{Z}/d\mathbb{Z}$, and the Galois action is given by
@@
For the curve $C\subset X$, we consider its image under the morphism $\rho$, denoted by $E=\rho(C)$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$.
@@
Since the action $\mathbb{Z}_{d}$ for the morphism $\rho$ is a Galois action, we see that $\rho^{-1}(E)$ is exactly the union of the curves in the $\mathbb{Z}_{d}$-orbit of $C$. It follows that $C_{i},i>1$ are all isomorphic to $C$ by automorphisms of $X$. Hence $C_{i}^{2}=C^{2}$ for all $i=1,\ldots,k$; and (i) follows.
Moreover, it is obvious that $g$ is a $\mathbb{Z}_{d}$-invariant polynomial in $(x,y,z,w)$. For each pair $i\neq j$, there exists an element $\alpha$ in the Galois group $\mathbb{Z}_{d}$ such that $\alpha\cdot C_{i}=C_{j}$, and then
$m_{i}=\mathrm{ord}_{C_{i}}(g)=\mathrm{ord}_{\alpha\cdot C_{i}}(\alpha\cdot g)=\mathrm{ord}_{C_{i}}(g)=m_{j}.$
Hence (ii) is true. ∎
3. Strategy of Proof
@@
By Proposition 2.1, there exists an irreducible polynomial $g(x,y,z)$ which cuts out a divisor on $X$ of the following form
@@
To investigate the Bounded Negativity Conjecture for $X_{d}$, we can safely assume that $k>1$; otherwise, $m^{2}C^{2}=d(\deg g)^{2}$ and thus $C^{2}\geq 0$. Therefore, we assume $k\geq 2$ in the sequel.
For convenience, denote by $D$ the curve defined by $x^{d}+y^{d}+z^{d}=0$ and $E$ the curve defined by $g=0$ in $\mathbb{P}^{2}$. Namely,
$D:x^{d}+y^{d}+z^{d}=0\qquad\mathrm{and}\qquad E:g(x,y,z)=0.$
Denote $e=\deg g=\deg E$.
Denote $\mathscr{C}$ the scheme on $X_{d}$ defined by $g=0$:

### v3

@@ Abstract.
We give an explicit formula for the self-intersection number of negative curves on Fermat surfaces. The formula offers us hints to prove the Bounded Negativity Conjecture for the Fermat surfaces by providing explicit negativity bounds, and it also offers the possibility of counterexamples to the conjecture.
@@ 1. Introduction
Let $\mathbb{P}^{n}$ be the complex projective space of dimension $n$. Let $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$ be the Fermat surface of degree $d$ in $\mathbb{P}^{3}$ with variables $x,y,z,w$, and let $D:x^{d}+y^{d}+z^{d}=0$ be the Fermat curve of degree $d$ in $\mathbb{P}^{2}$ with variables $x,y,z$. We shall study self-intersection number of negative curves on $X_{d}$. The main objective in this paper is to establish the following theorem.
@@
Let $d\geq 5$. Suppose that $C$ is a reduced and irreducible curve on the Fermat surface $X_{d}$ with self-intersection number $C^{2}<2-d<0$. Then the following properties hold.
  1. (1)
    There exists a unique reduced and irreducible curve $E\subset\mathbb{P}^{2}$ defined by a degree $e$ polynomial $g(x,y,z)$, such that the cone in $\mathbb{P}^{3}$ determined by $g$ contains $C$.
  2. (2)
    There exists $k\geq 2$ and $k|d$ such that
@@
    (1.1) $+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E).$
    In particular, when $d$ is prime, we have $k=d$ and
@@
    In the above formulae, $\mu_{p}(E)$ denotes the Milnor number of the plane curve $E$ at the point $p$, and $\mu_{[x,y,z,0]}(C)$ denotes the Milnor number of $C$ at $[x,y,z,0]$, and $(D\cap E)$ the number of points in the set $D\cap E$.
The technical condition $C^{2}<2-d$ in Theorem 1.1 is given to exclude exceptional cases where $C^{2}$ can be calculated explicitly. This ensures a concise formulation while maintaining generality. The exceptional cases are investigated in Section 3.
The computation of self-intersection number of negative curves, as in Theorem 1.1, is closely connected with an old folklore conjecture among experts in algebraic surface theory, commonly known as the Bounded Negativity Conjecture. We formulate this conjecture as follows.
@@
For any given smooth complex projective surface $X$, there exists a number $B_{X}\geq 0$ which depends only on $X$ such that for any reduced curve $C$ on $X$, the self-intersection number $C^{2}$ is bounded below by $B_{X}$, i.e., $C^{2}\geq-B_{X}$.
The origins of this conjecture are unclear, but it can at least be traced back to the Italian mathematician Federigo Enriques (1871-1946), and it has been an open problem for one hundred years. For a more comprehensive overview of the history, we refer to [3, p.1878].
@@
For any given smooth complex projective surface $X$, there exists a number $b_{X}\geq 0$ which depends only on $X$ such that for any reduced and irreducible curve $C$ on $X$, the self-intersection number $C^{2}$ is bounded below by $b_{X}$, i.e., $C^{2}\geq-b_{X}$.
@@
The result in Theorem 1.1 implies that under some conditions, the Bounded Negativity Conjecture is true for the Fermat surface $X_{d}$.
@@
    for some constant $\lambda>0$ which may depend on $d$; or
  2. (b)
    $\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}<1;$
where, in the above two formulas, the curve $E$ ranges over all irreducible projective plane curves of degree $e$.
On the other hand, we hope that the formula for $C^{2}$ in Theorem 1.1 also provides the possibility of counterexamples to the Bounded Negativity Conjecture on Fermat surfaces, although no explicit counterexamples have been constructed.
In the sequel, we will only investigate the self-intersection number of reduced and irreducible curves. For convenience, we denote $C$ such a curve on a given smooth projective surface $X$.
It is evident that the Bounded Negativity Conjecture holds for $X$ with $K_{X}\geq 0$, as can be seen by applying the adjunction formula:
@@
In particular, for $d\leq 4$, the bounded negativity of $X_{d}$ follows from this argument, since the canonical divisor of $X_{d}$ is $K_{X_{d}}=(d-4)H$ where $H$ is the hyperplane section. However, for $d\geq 5$, the canonical divisor $K_{X_{d}}$ is ample and $X_{d}$ is of general type. To the author’s knowledge, there are currently no effective methods available to deal with the Bounded Negativity of such surfaces, except for the weak Bounded Negativity Conjecture proved in [12].
Another natural approach to establish the Bounded Negativity for a surface is to show that there are only finitely many negative curves on it. However, it is proved in [3, Thm.4.1] that for any $m>0$, there exists a smooth projective surface which admits infinitely many smooth irreducible curves of self-intersection number m-m. For the Fermat surface $X_{d}$, we do not know whether the number of negative curves on $X_{d}$ is finite or not.
Other strategies to prove the Bounded Negativity Conjecture include investigating the non-isomorphic surjective endomorphisms or establishing dominance by other surfaces that have bounded negativity [3, Prop.2.1], exploring the birational invariance of bounded negativity and the $H$-constant [2], using the equivalence between bounded negativity and the boundedness of the denominators in the Zariski decomposition [4], discussing the weighted Bounded Negativity Conjecture [13, 9, 10], and studying the connections between the Bounded Negativity Conjecture and other unsolved problems, such as SHGH conjecture or the bounded cohomology conjecture [1, 7].
However, despite providing invaluable insights in understanding the Bounded Negativity Conjecture, these methods either deal with only very special cases, e.g., rational surfaces or K3 surfaces, or just offer general frameworks that reduce the conjecture to other open questions. They rely on general formulae, such as the adjunction formula and Riemann-Roch formula, to establish bounds for the intersection number $C^{2}$ of a reduced and irreducible curve $C$, without using specific information about $C$, such as its equation. In addition, the negativity bound, even if exists, remains implicit in most cases.
Can we estimate $C^{2}$ from the equation of the curve $C$? Can we give a concrete example of a surface with the Bounded Negativity by performing explicit computation through the defining equation of the surface, thereby yielding an explicit negativity bound?
@@
Suppose $X$ and $Y$ are two smooth complex projective surfaces belonging to a single family of complex projective surfaces, then the Bounded Negativity Conjecture is true for $X$ if and only if it is true for $Y$.
Assuming the conjectures 1.5,1.6 above, we are motivated to prove the validity of the Bounded Negativity Conjecture for a specific surface in $\mathbb{P}^{3}$. Theorem 1.1 is proved for this purpose.
Note that a general surface in $\mathbb{P}^{3}$ has Picard number 1, which follows from the Noether-Lefschetz theorem; hence for a general surface $X\subset\mathbb{P}^{3}$, we may take $b_{X}=0$. Therefore, the Bounded Negativity Conjecture for surfaces in $\mathbb{P}^{3}$ follows from Conjecture 1.6 without explicit computation. However, the Fermat surface does not have Picard number 1, and our computation in Theorem 1.1 is of independent interest, particularly in providing a nontrivial explicit bound $c_{d}$ in Conjecture 1.5.
The formulae in Theorem 1.1 show how the Bounded Negativity Conjecture is connected with the singularity theory of curves. The key term in formula (2) is the sum $\sum\limits_{p\in D\cap E}\mu_{p}(E)$. It is notable that this sum is taken over the intersection points of $D$ and $E$, not all the singular points of $E$.
To conclude the introduction, we give an outline of the proof of Theorem 1.1. We represent the Fermat surface $X_{d}$ as a branched Galois cover of $\mathbb{P}^{2}$:
$\rho:\ X_{d}\to\mathbb{P}^{2},\quad[x,y,z,w]\mapsto[x,y,z].$
Given a negative curve $C$ on $X_{d}$, let $E=\rho(C)$ and $\mathscr{C}=\rho^{*}E$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$, whose defining equation is denoted by $g=g(x,y,z)$. It is clear that $g$ can be seen as a section of the line bundle $\mathscr{O}_{X_{d}}\left((\deg g)H\right)$ which vanishes on $C$. Then we prove that $\mathscr{C}$ is a reduced curve on $X_{d}$; see Proposition 4.4. Moreover, $\mathscr{C}$ is of the following form
@@
As a result, to compute $C^{2}$, we need to calculate the sum of intersection numbers on the right hand side. Finally, the intersection numbers can be calculated in a local way at any intersection point. Locally, the curve $\mathscr{C}$ can be represented as the curve in $(\mathbb{C}^{3},0)$ defined by
@@
The author would like to thank the organizers of International Symposium on Singularities and Applications, Dec. 9–13, 2024 in Sanya, China. The stimulating atmosphere of the conference inspired him to conceive the idea of performing computation using Puiseux series. The author also thanks Professor Feng Hao for helpful discussions.
@@ 2. Curves on surfaces of Sebastiani-Thom type
If $X\subset\mathbb{P}^{3}$ is a smooth surface defined by the following equation:
@@
where $f$ is a form of degree $d$, then $X$ is said to be of Sebastiani-Thom type. The Fermat surface $X_{d}$ is a typical example.
We show that a curve on $X$ is “almost” defined by a homogeneous polynomial in the following result.
@@
Let $X:f(x,y,z)+w^{d}=0$ be a smooth surface in $\mathbb{P}^{3}$ and $C$ a reduced and irreducible curve on $X$. Then there exists a unique irreducible form $g(x,y,z)$ up to a multiplicative constant, such that the divisor on $X$ cut out by $g$ can be written as
@@
where $C_{1}=C,C_{2},\ldots,C_{k}$ are distinct reduced and irreducible curves and $m\geq 1$ is an integer, and, in addition, the self-intersection numbers $C_{i}^{2},i=1,\ldots,k$ are all equal to $C^{2}$.
@@
It is a branched Galois covering map with Galois group $\mathbb{Z}_{d}=\mathbb{Z}/d\mathbb{Z}$, the Galois action given by
@@
Denote $E=\rho(C)$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$.
@@
Since the action $\mathbb{Z}_{d}$ for the morphism $\rho$ is a Galois action, $\rho^{-1}(E)$ is exactly the union of the curves in the $\mathbb{Z}_{d}$-orbit of $C$. It follows that for each pair $i\neq j$, there exists an element $\alpha$ in the Galois group $\mathbb{Z}_{d}$ such that $\alpha\cdot C_{i}=C_{j}$. Now it is obvious that $g$ is a $\mathbb{Z}_{d}$-invariant polynomial in $(x,y,z,w)$; thus,
$m_{i}=\mathrm{ord}_{C_{i}}(g)=\mathrm{ord}_{\alpha\cdot C_{i}}(\alpha\cdot g)=\mathrm{ord}_{C_{j}}(g)=m_{j}.$
Hence $m_{1}=\cdots=m_{k}=m$ for some $m\geq 1$. In addition, since $C_{i},i>1$ are all isomorphic to $C$ by automorphisms of $X$, we have $C_{i}^{2}=C^{2}$ for all $i=1,\ldots,k$.
It remains to show that the polynomial $g$ is unique up to a constant factor. Indeed, $g$ must be the defining polynomial of the irreducible curve $E$ since $g$ depends only the variables $x,y,z$, without involving the variable $w$. ∎
3. Strategy of Proof of Theorem 1.1
@@
By Proposition 2.1, there exists an irreducible polynomial $g(x,y,z)$ which cuts out a divisor on $X_{d}$ of the following form
@@
If $k=1$, we have $m^{2}C^{2}=d(\deg g)^{2}$ and thus $C^{2}\geq 0$. Therefore, to investigate the Bounded Negativity Conjecture for $X_{d}$, we assume $k\geq 2$ in the sequel.
Recall that $D:x^{d}+y^{d}+z^{d}=0$ is the Fermat curve. Let $E:g(x,y,z)=0$ be the plane curve defined by $g$ in $\mathbb{P}^{2}$, and let $e=\deg g=\deg E$.
Furthermore, denote $\mathscr{C}$ the scheme on $X_{d}$ defined by $g=0$: